# MiLoRA: Harnessing Minor Singular Components for Parameter-Efficient LLM Finetuning

### v1

@@ Abstract
Efficient finetuning of large language models (LLMs) aims to adapt the LLMs with reduced computation and memory cost. Previous LoRA-based approaches initialize the low-rank matrices with gaussian distribution and zero values, while keeping the original weight matrices frozen. However, the trainable model parameters optimized in an unguided subspace might have interference with the well-learned subspace of the pretrained weight matrix. In this paper, we propose MiLoRA, a simple yet effective LLM finetuning approach that only updates the minor singular components of the weight matrix while keeping the principle singular components frozen. It is observed that the minor matrix corresponds to the noisy or long-tail information, while the principle matrix contains important knowledge. The MiLoRA initializes the low-rank matrices within a subspace that is orthogonal to the principle matrix, thus the pretrained knowledge is expected to be well preserved. During finetuning, MiLoRA makes the most use of the less-optimized subspace for learning the finetuning dataset. Extensive experiments on commonsense reasoning, math reasoning and instruction following benchmarks present the superior performance of our method.
@@ 1 Introduction
Large language models (Brown et al., 2020; Ouyang et al., 2022; Touvron et al., 2023; Jiang et al., 2023, LLMs) have demonstrated superior performance on various tasks (Zheng et al., 2023), such as math reasoning (Wang et al., 2024) and question answering (Ivison et al., 2023). These models are pretrained with the next token prediction task (Brown et al., 2020) on large web-scale data, then finetuned with instruction data as well as human preference data (Ouyang et al., 2022; Yu et al., 2023; Cui et al., 2023) for different downstream tasks. Fully finetuning is commonly employed to unlock the complete potential of LLMs, however, optimizing all model parameters necessitates substantial and restrictive computational resources (Touvron et al., 2023; Jiang et al., 2023), which hampers the utilization of LLMs across diverse scenarios.
@@
Parameter-efficient finetuning (Hu et al., 2021, 2023b; Liu et al., 2022; Zhang et al., 2024; Liu et al., 2024a, PEFT) aims at reducing the computational and GPU memory cost for finetuning of pretrained models. Low-rank adaptation (Hu et al., 2021, LoRA) is one of the most widely used PEFT methods for LLM finetuning. It assumes the change of linear model weights to be low-rank (Li et al., 2018; Aghajanyan et al., 2021). For each selected weight matrix, it only updates two low-rank matrices while keeping the pretrained weight frozen. During inference, the low-rank matrices are merged into pretrained linear weights, thus no additional computation or memory cost is introduced. Recently, researchers explore different LoRA-based variants for efficient LLM finetuning (Zhang et al., 2023; Pan et al., 2024; Kopiczko et al., 2024; Liu et al., 2024a; Meng et al., 2024). However, most existing LoRA-based works randomly initialize the low-rank matrices and optimize the trainable parameters in an unguided subspace. We suspect this strategy may override important pretrained features, thus influencing the performance of low-rank adaptation methods (Dou et al., 2024).
@@
In this paper, we propose Minor singular component based Low Rank Adaptation (MiLoRA) for efficient LLM finetuning. MiLoRA has a similar framework as vanilla LoRA, but employs a different initialization schedule. Specifically, a weight matrix ${\bm{W}}$ is decomposed with singular value decomposition (SVD) algorithm. Based on the magnitude of the singular values, we divide ${\bm{W}}$ into two components: the principle matrix ${\bm{W}}_{p}$ corresponding to large singular values and the minor matrix ${\bm{W}}_{m}$ corresponding to small singular values. We argue that the principle matrix captures the essence of the pretrained knowledge, whereas the minor matrix is suboptimal with noisy or long-tail information. It is supported by previous works (Hajimolahoseini et al., 2021; Sharma et al., 2024) that using principle low-rank approximation matrices can achieve comparable or even better performance to full parameters.
@@
Motivated by these observations, we keep the principle matrix ${\bm{W}}_{p}$ frozen and adapt the minor singular components during finetuning. The low-rank matrices ${\bm{A}}$ and ${\bm{B}}$ in LoRA framework are initialized with the minor matrix ${\bm{W}}_{m}$. Since the trainable low-rank matrices are initialized in a subspace orthogonal to the principle matrix, MiLoRA is expected to effectively learn from finetuning tasks while better preserving and utilizing the pretrained knowledge. To maintain the capability of the pretrained model at the start of finetuning, vanilla LoRA explicitly initializes ${\bm{B}}$ with zeros. In contrast, MiLoRA naturally satisfies this requirement, as the pretrained weight matrix ${\bm{W}}$ equals to the frozen principle part ${\bm{W}}_{p}$ plus the low-rank part ${\bm{W}}_{m}={\bm{B}}{\bm{A}}$. We conduct extensive experiments on commonsense reasoning, math reasoning and instruction following benchmarks. The experimental results show that MiLoRA consistently outperforms LoRA without sacrificing training or inference efficiency, such as commonsense reasoning (+1.6/+1.1 on LLaMA2-7B/LLaMA3-8B), math reasoning (+1.92 on LLaMA2-7B), and instruction following (+1.4 on LLaMA2-7B). [^1] Our code and model will be made public.
@@ 2 Preliminaries
Singualr Value Decomposition
Given a matrix ${\bm{W}}\in\mathbb{R}^{m\times n}$, its singular value decomposition (SVD) is denoted as ${\bm{W}}={\bm{U}}\bm{\Sigma}{\bm{V}}^{\top}$, where ${\bm{U}}=\left[u_{1},u_{2},\cdots,u_{m}\right]\in\mathbb{R}^{m\times m}$, ${\bm{V}}=\left[v_{1},v_{2},\cdots,v_{n}\right]\in\mathbb{R}^{n\times n}$. The columns of ${\bm{U}}$ are the left singular vectors, and the columns of ${\bm{V}}$ are the right singular vectors. The $\bm{\Sigma}\in\mathbb{R}^{m\times n}$ is a diagonal matrix containing the singular values of ${\bm{W}}$ in descending order. Without loss of generality, we suppose $m\leq n$ to simplify the notation. The SVD of ${\bm{W}}$ can be reformulated as in Equation 1.
@@
where ${\bm{W}}^{(0)}$ and $\Delta{\bm{W}}$ refer to the pretrained weight and weight change, respectively. The $\alpha$ and $r$ are hyperparameters of scaling-factor and LoRA rank ($r\ll\min(m,n)$). During finetuning, the pretrained matrix ${\bm{W}}^{(0)}$ is kept frozen. It significantly diminishes the number of trainable parameters as both ${\bm{A}}$ and ${\bm{B}}$ matrices are low-rank. The ${\bm{B}}$ matrix is initialized with zero while ${\bm{A}}$ matrix adopts a random gaussian distribution with zero mean value. This initialization strategy ensures the $\Delta{\bm{W}}=0$ at the beginning of training. The LoRA method only modifies the linear matrices in the Transformer model. The low-rank matrices can be easily merged into the pretrained linear matrix to get updated for inference, which does not require additional computing and GPU memory compared with fully finetuning. However, vanilla LoRA method fails to select the optimal subspace for updating the model parameters, as the low-rank matrices ${\bm{A}}$ and ${\bm{B}}$ are randomly initialized. This might potentially detract from the pretrained knowledge encoded in the principal weight matrix.
Figure 1: (a) MiLoRA method splits the pretrained linear weight matrix into two parts, the minor singular component (${\bm{U}}_{m},{\bm{V}}_{m},{\bm{\Sigma}}_{m}$) is used to initialize the low-rank matrices ${\bm{A}}_{m}$ and ${\bm{B}}_{m}$ in the LoRA framework. (b) During finetuning, only the low-rank matrices ${\bm{A}}_{m}$ and ${\bm{B}}_{m}$ are updated while the principle matrix ${\bm{W}}_{p}$ is frozen, which is similar as the practice of LoRA method.
@@ 3 Methodology
The proposed MiLoRA is a simple yet effective PEFT approach that selects the minor singular components of the weight matrices for optimization. As observed in LASER (Sharma et al., 2024), the minor singular components of weight matrices contain noisy or long-tail information, while the principle singular components are responsible for important features across tasks. Therefore, MiLoRA is designed to effectively learn from the finetuning dataset by adapting minor singular components, while maintaining the knowledge encoded in the pretrained model. On the contrary, vanilla LoRA fails to constrain the optimization subspace as its low-rank matrices are randomly initialized. Specifically, suppose $m\leq n$, MiLoRA splits each linear weight matrix into two matrices according to the corresponding singular values: principal matrix ${\bm{W}}_{p}$ and minor matrix ${\bm{W}}_{m}$ (see Figure 1):
${\bm{W}}={\bm{W}}_{p}+{\bm{W}}_{m}=\sum_{i=1}^{m-r}\sigma_{i}u_{i}v_{i}^{\top}+\sum_{i=m-r+1}^{m}\sigma_{i}u_{i}v_{i}^{\top},$ (3)
@@
The decomposition can be also reformulated in a matrix form. The ${\bm{U}}$ matrix in the SVD of ${\bm{W}}$ can be reformulated as ${\bm{U}}=\left[{\bm{U}}_{p},{\bm{U}}_{m}\right]$, where ${\bm{U}}_{p}=\left[u_{1},u_{2},\cdots,u_{m-r}\right]$ and ${\bm{U}}_{m}=\left[u_{m-r+1},u_{m-r+2},\cdots,u_{m}\right]$ are left singular vectors corresponding to principle and minor singular values, respectively. The ${\bm{V}}$ and ${\bm{\Sigma}}$ matrices can be reformulated in a similar way. The SVD of ${\bm{W}}$ can be expressed as:
${\bm{W}}={\bm{U}}{\bm{\Sigma}}{\bm{V}}^{\top}={\bm{U}}_{p}{\bm{\Sigma}}_{p}{\bm{V}}_{p}^{\top}+{\bm{U}}_{m}{\bm{\Sigma}}_{m}{\bm{V}}_{m}^{\top}={\bm{W}}_{p}+{\bm{W}}_{m}$ (4)
@@
${\bm{W}}_{m}={\bm{U}}_{m}{\bm{\Sigma}}_{m}{\bm{V}}_{m}^{\top}=({\bm{U}}_{m}\sqrt{{\bm{\Sigma}}_{m}})(\sqrt{{\bm{\Sigma}}_{m}}{\bm{V}}_{m}^{\top})={\bm{B}}_{m}{\bm{A}}_{m}$ (5)
This strategy has two benefits: 1) It encourages the model to learn in the less-optimized subspace spanned by the minor singular vectors, thus reducing the interference with the well-learned pretrained knowledge encoded in the principle singualr components. 2) Unlike vanilla LoRA which requires tuning the scaling factor $\alpha$ and the initialization hyperparameters, ours does not require any hyperparameter except the rank $r$. To make our method simple and training-efficient, we do not incorporate the regularization loss which imposes strict constraints on the orthogonality between ${\bm{W}}_{m}$ and ${\bm{W}}_{p}$. Our experiment results show that our scheme works quite well in balancing learning from finetuning dataset and preverving pretrained knowledge.
@@ 4 Experiments
To evaluate our method against other PEFTs, we conduct experiments across three diverse tasks covering more than 10 datasets. Our objective is to illustrate the effectiveness of our method across various scenarios. We benchmark against LoRA (Hu et al., 2021) and PiSSA (Meng et al., 2024).
  • PiSSA (Meng et al., 2024) shares the same framework as LoRA, but it employs a more sophisticated initialization approach. Specifically, they initialize ${\bm{A}}$ and ${\bm{B}}$ with principal singular values and singular vectors of the pre-trained weight ${\bm{W}}$. Given that the principal singular values and vectors capture the essence of a matrix, PiSSA is hoped to better approximate the outcomes of full-parameter finetuning by changing the essential parts while freezing the “noisy” parts.
The MiLoRA method optimizes the low-rank matrices within a less-optimized subspace, which can be combined with other LoRA-based PEFT methods (Liu et al., 2024a; Zhang et al., 2023; Pan et al., 2024). We leave the exploration of combination MiLoRA with LoRA variants as future work.
4.1 Commonsense reasoning
Setup
We finetune LLaMA2-7B (Touvron et al., 2023) and LLaMA3-7B (AI@Meta, 2024) on Commonsense170K (Hu et al., 2023a). Eight commonsense reasoning datasets are used for evaluation, including BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-e, ARC-c (Clark et al., 2018), and OBQA (Mihaylov et al., 2018). The task is formulated as a multiple-choice problem. The best checkpoint is chosen by their validation set loss. We report accuracy (%) for all datasets. We use the implementation of LLM-Adapters (Hu et al., 2023b) [^2] https://github.com/AGI-Edgerunners/LLM-Adapters.
Hyperparameter
We use the same hyperparameter configurations as Hu et al. (2023a) without tuning for all methods. Details can be found in Table 8 of Appendix A.2. We denote this hyperparameter setup as LLM-Adapters.
Results
Table 1: Commonsense reasoning evaluation results for LLaMA2-7B and LLaMA3-8B on eight tasks. Performance results of baseline methods are taken from Liu et al. (2024a). Bold numbers indicate the highest performance scores for each dataset across the different PEFT methods for the corresponding model.
@@ Results
We report results in Table 1. We also include the ChatGPT baseline reported in Liu et al. (2024a) as a reference, which is obtained with gpt-3.5-turbo API using a zero-shot Chain of Thought (Wei et al., 2022). MiLoRA consistently surpasses all baseline methods across both LLaMA2-7B and LLaMA3-8B models. On LLaMA2-7B, MiLoRA outperforms LoRA and PiSSA by 1.6 and 5.4 average accuracy scores. On LLaMA3-8B, MiLoRA exceeds LoRA and PiSSA by 1.1 and 6.5 average accuracy scores. These outcomes suggest the effectiveness of MiLoRA as a PEFT method.
4.2 Math reasoning
Setup
We finetune LLaMA2-7B (Touvron et al., 2023) on the MetaMathQA dataset (Yu et al., 2023), which contains 395K samples augmented from the training set of GSM8K (Cobbe et al., 2021) and MATH(Hendrycks et al., 2021). We use test set of GSM8K and MATH for evaluation and report results on the last checkpoint. For each test set, we report the exact match ratio against the ground truth. We use the implementation of PiSSA(Meng et al., 2024) [^3] https://github.com/GraphPKU/PiSSA.
Hyperparameter
As MetaMathQA has a large training dataset, we set the rank to 64. For all other hyperparameters, we follow the llm-adapter hyperparameter setup shown in Table 8 of Appendix A.2. We limit all samples to a maximum length of 2048 tokens.
Results
@@
LoRA 60.58 16.88 38.73
PiSSA 58.23 15.84 37.04
MiLoRA 63.53 17.76 40.65
We report results in Table 2. We also include the full finetuning results from (Yu et al., 2023) as a reference. We can see that MiLoRA achieves the best performance on LLaMA2-7B. However, it still underperforms full finetuning by an average of 2.55 EM score. This suggests that PEFT methods still have room for improvement to fully match the performance of full finetuning.
4.3 Instruction-following
Table 3: Instruction tuning evaluation results for instruction-tuned LLaMA2 7B with Alpaca-Eval v1.0. Performance results of baseline methods are taken from Ivison et al. (2023).

### v2

@@ Abstract
Efficient finetuning of large language models (LLMs) aims to adapt the LLMs with reduced computational and memory cost. Previous LoRA-based approaches initialize the low-rank matrices with Gaussian distribution and zero values while keeping the original weight matrices frozen. However, the trainable model parameters optimized in an unguided subspace might interfere with the well-learned subspace of the pretrained weight matrices. In this paper, we propose MiLoRA, a simple yet effective LLM finetuning approach that only updates the minor singular components of the weight matrix while keeping the principal singular components frozen. It is observed that the minor matrix corresponds to the noisy or long-tail information, while the principal matrix contains important knowledge. The MiLoRA initializes the low-rank matrices within a subspace that is orthogonal to the principal matrix, thus the pretrained knowledge is expected to be well preserved. During finetuning, MiLoRA makes the most use of the less-optimized subspace for learning the labeled dataset. Extensive experiments on commonsense reasoning, math reasoning, instruction following and visual instruction following benchmarks present the superior performance of our method.
MiLoRA: Harnessing Minor Singular Components for Parameter-Efficient LLM Finetuning
Hanqing Wang1,  Yixia Li2,  Shuo Wang3,  Guanhua Chen2,  Yun Chen1 1Shanghai University of Finance and Economics 2Southern University of Science and Technology   3Tsinghua University
@@ 1 Introduction
Large language models (Brown et al., 2020; Ouyang et al., 2022; Touvron et al., 2023; Jiang et al., 2023, LLMs) have demonstrated superior performance on various tasks (Zheng et al., 2023), such as math reasoning (Wang et al., 2024a) and question answering (Ivison et al., 2023). These models are pretrained with the next token prediction task (Brown et al., 2020) on large web-scale data, then finetuned with instruction data as well as human preference data (Ouyang et al., 2022; Yu et al., 2023; Cui et al., 2023) for different downstream tasks. Fully finetuning is commonly employed to unlock the complete potential of LLMs, however, optimizing all model parameters necessitates substantial and restrictive computational resources (Touvron et al., 2023; Jiang et al., 2023), which hampers the utilization of LLMs across diverse scenarios.
@@
Parameter-efficient finetuning (Hu et al., 2021, 2023; Liu et al., 2022; Zhang et al., 2024; Liu et al., 2024b, PEFT) aims at reducing the computational and GPU memory cost for finetuning of pretrained models. Low-rank adaptation (Hu et al., 2021, LoRA) is one of the most widely used PEFT methods for LLM finetuning. It assumes the change of linear model weights to be low-rank (Li et al., 2018; Aghajanyan et al., 2021). For each selected weight matrix, it only updates two low-rank matrices while keeping the pretrained weight frozen. During inference, the low-rank matrices are merged into the pretrained linear weights, thus no additional computational or memory cost is introduced. Recently, researchers have explored different LoRA-based variants for efficient LLM finetuning (Zhang et al., 2023; Pan et al., 2024; Kopiczko et al., 2024; Liu et al., 2024b; Meng et al., 2024). However, most existing LoRA-based methods randomly initialize the low-rank matrices and optimize the trainable parameters in an unguided subspace. We suspect this strategy may override important pretrained features, thus degrading the performance of low-rank adaptation methods (Dou et al., 2024).
@@
In this paper, we propose Minor singular component based Low Rank Adaptation (MiLoRA) for efficient LLM finetuning. MiLoRA has a similar framework as LoRA but employs a different initialization schedule. Specifically, a weight matrix ${\bm{W}}$ is decomposed with the singular value decomposition (SVD) algorithm. Based on the magnitude of the singular values, we divide ${\bm{W}}$ into two components: the principal matrix ${\bm{W}}_{p}$ corresponding to large singular values and the minor matrix ${\bm{W}}_{m}$ corresponding to small singular values. We argue that the principal matrix captures the essence of the pretrained knowledge, whereas the minor matrix is suboptimal with noisy or long-tail information. It is supported by previous works (Hajimolahoseini et al., 2021; Sharma et al., 2024) that principal low-rank approximation can achieve comparable or even better performance to full parameters.
@@
Motivated by these observations, we freeze the principal matrix ${\bm{W}}_{p}$ and adapt the minor singular components during finetuning. The low-rank matrices ${\bm{A}}$ and ${\bm{B}}$ in LoRA framework are initialized with the minor matrix ${\bm{W}}_{m}$. Since the trainable low-rank matrices are initialized in a subspace orthogonal to the principal matrix, MiLoRA is expected to effectively learn from finetuning tasks while better preserving and utilizing the pretrained knowledge. To maintain the capability of the pretrained model at the start of finetuning, vanilla LoRA explicitly initializes ${\bm{B}}$ with zeros. In contrast, MiLoRA naturally satisfies this requirement, as the pretrained weight matrix ${\bm{W}}$ equals the frozen principal part ${\bm{W}}_{p}$ plus the low-rank part ${\bm{W}}_{m}={\bm{B}}{\bm{A}}$. We conduct extensive experiments on commonsense reasoning, math reasoning, instruction following and visual instruction following benchmarks. The experimental results show that MiLoRA consistently outperforms LoRA without sacrificing training or inference efficiency, such as commonsense reasoning (+1.6/+1.1 on LLaMA2-7B/LLaMA3-8B), math reasoning (+2.0 on LLaMA2-7B), instruction following (+1.4 on LLaMA2-7B) and visual instruction following (+1.4 on LLaVA1.5-7B). [^1] Our code and model will be made public.
@@ 2 Preliminaries
Singular Value Decomposition
Given a matrix ${\bm{W}}\in\mathbb{R}^{m\times n}$, its singular value decomposition is denoted as ${\bm{W}}={\bm{U}}\bm{\Sigma}{\bm{V}}^{\top}$, where ${\bm{U}}=\left[u_{1},u_{2},\cdots,u_{m}\right]\in\mathbb{R}^{m\times m}$, ${\bm{V}}=\left[v_{1},v_{2},\cdots,v_{n}\right]\in\mathbb{R}^{n\times n}$. The columns of ${\bm{U}}$ are the left singular vectors, and the columns of ${\bm{V}}$ are the right singular vectors. The $\bm{\Sigma}\in\mathbb{R}^{m\times n}$ is a diagonal matrix containing the singular values of ${\bm{W}}$ in descending order.
Without loss of generality, we suppose $m\leq n$ to simplify the notation. The SVD of ${\bm{W}}$ can be reformulated as in Equation 1.
@@
where ${\bm{W}}^{(0)}$ and $\Delta{\bm{W}}$ refer to the pretrained weight and weight change, respectively. The $\alpha$ and $r$ are hyperparameters of scaling-factor and LoRA rank ($r\ll\min(m,n)$). During finetuning, the pretrained matrix ${\bm{W}}^{(0)}$ is kept frozen. It significantly diminishes the number of trainable parameters as both ${\bm{A}}$ and ${\bm{B}}$ matrices are low-rank. The ${\bm{B}}$ matrix is initialized with zero while ${\bm{A}}$ matrix adopts a random Gaussian distribution with zero mean value.
This initialization strategy ensures the $\Delta{\bm{W}}=0$ at the beginning of training. The LoRA method only modifies the linear matrices in the Transformer model. The low-rank matrices can be easily merged into the pretrained linear matrix to get updated for inference, which does not require additional computing and GPU memory compared with full finetuning. However, the vanilla LoRA method fails to select the optimal subspace for updating the model parameters, as the low-rank matrices ${\bm{A}}$ and ${\bm{B}}$ are randomly initialized. This might potentially detract from the pretrained knowledge encoded in the pretrained weight matrix.
Figure 1: (a) MiLoRA method splits the pretrained linear weight matrix into two parts, the minor singular component (${\bm{U}}_{m},{\bm{V}}_{m},{\bm{\Sigma}}_{m}$) is used to initialize the low-rank matrices ${\bm{A}}_{m}$ and ${\bm{B}}_{m}$ in the LoRA framework. (b) During finetuning, only the low-rank matrices ${\bm{A}}_{m}$ and ${\bm{B}}_{m}$ are updated while the principal matrix ${\bm{W}}_{p}$ is frozen, which is similar to the practice of LoRA method.
@@ 3 Methodology
The proposed MiLoRA is a simple yet effective PEFT approach that selects the minor singular components of the weight matrices for optimization. As observed in LASER (Sharma et al., 2024), the minor singular components of weight matrices contain noisy or long-tail information, while the principal singular components are responsible for important features across tasks. Therefore, MiLoRA is designed to effectively learn from the finetuning dataset by adapting minor singular components, while maintaining the knowledge encoded in the pretrained model. On the contrary, vanilla LoRA fails to constrain the optimization subspace as its low-rank matrices are randomly initialized. Specifically, suppose $m\leq n$, MiLoRA splits each linear weight matrix into two matrices according to the corresponding singular values: principal matrix ${\bm{W}}_{p}$ and minor matrix ${\bm{W}}_{m}$ (see Figure 1):
${\bm{W}}$ $={\bm{W}}_{p}+{\bm{W}}_{m}$ (3)
$=\sum_{i=1}^{m-r}\sigma_{i}u_{i}v_{i}^{\top}+\sum_{i=m-r+1}^{m}\sigma_{i}u_{i}v_{i}^{\top},$
@@
The decomposition can be also reformulated in a matrix form. The ${\bm{U}}$ matrix in the SVD of ${\bm{W}}$ can be reformulated as ${\bm{U}}=\left[{\bm{U}}_{p},{\bm{U}}_{m}\right]$, where ${\bm{U}}_{p}=\left[u_{1},u_{2},\cdots,u_{m-r}\right]$ and ${\bm{U}}_{m}=\left[u_{m-r+1},u_{m-r+2},\cdots,u_{m}\right]$ are left singular vectors corresponding to principal and minor singular values, respectively. The ${\bm{V}}$ and ${\bm{\Sigma}}$ matrices can be reformulated similarly. The SVD of ${\bm{W}}$ can be expressed as:
${\bm{W}}$ $={\bm{U}}{\bm{\Sigma}}{\bm{V}}^{\top}={\bm{U}}_{p}{\bm{\Sigma}}_{p}{\bm{V}}_{p}^{\top}+{\bm{U}}_{m}{\bm{\Sigma}}_{m}{\bm{V}}_{m}^{\top}$ (4)
$={\bm{W}}_{p}+{\bm{W}}_{m}$
@@
${\bm{W}}_{m}$ $={\bm{U}}_{m}{\bm{\Sigma}}_{m}{\bm{V}}_{m}^{\top}$ (5)
$=({\bm{U}}_{m}\sqrt{{\bm{\Sigma}}_{m}})(\sqrt{{\bm{\Sigma}}_{m}}{\bm{V}}_{m}^{\top})={\bm{B}}_{m}{\bm{A}}_{m},$
This strategy has two benefits: 1) It encourages the model to learn in the less-optimized subspace spanned by the minor singular vectors, thus reducing the interference with the well-learned pretrained knowledge encoded in the principal singular components. 2) Unlike vanilla LoRA which requires tuning the scaling factor $\alpha$ and the initialization hyperparameters, ours does not require any hyperparameter except the rank $r$.
To make our method simple and training-efficient, we do not incorporate the regularization loss which imposes strict constraints on the orthogonality between ${\bm{W}}_{m}$ and ${\bm{W}}_{p}$. Our experiment results show that our scheme works quite well in balancing learning from the finetuning dataset and preserving pretrained knowledge.
@@ 4 Experiments
Table 1: Commonsense reasoning evaluation results for LLaMA2-7B and LLaMA3-8B on eight tasks. Results are cited from Liu et al. (2024b). Bold numbers indicate the highest performance scores for each dataset across the different PEFT methods for the corresponding model.
@@
LoRA 60.6 16.9 38.7
PiSSA 58.2 15.8 37.0
MiLoRA 63.5 17.8 40.7
To evaluate MiLoRA against other PEFTs, we conduct experiments on both NLP and multimodal tasks covering 18 datasets. We benchmark against LoRA (Hu et al., 2021) and PiSSA (Meng et al., 2024). All of our experiments are conducted on 8 NVIDIA L40 GPUs unless otherwise specified.
  • PiSSA (Meng et al., 2024) shares the same framework as LoRA, but employs a more sophisticated initialization approach. They initialize ${\bm{A}}$ and ${\bm{B}}$ with principal singular values and singular vectors of the pre-trained weight ${\bm{W}}$. Given that the principal components capture the essence of a matrix, PiSSA is expected to better approximate full finetuning by changing the essential parts while freezing the “noisy” parts.
4.1 Experiments on Large Language Model
Models and Datasets
We evaluate and compare MiLoRA with baselines on three different types of LLM downstream tasks.
1. Commonsense reasoning: We finetune LLaMA2-7B (Touvron et al., 2023) and LLaMA3-7B (AI@Meta, 2024) on Commonsense170K (Hu et al., 2023). Eight commonsense reasoning datasets are used for evaluation, including BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-e, ARC-c (Clark et al., 2018), and OBQA (Mihaylov et al., 2018). The task is formulated as a multiple-choice problem. We report accuracy (%) for all datasets on the best checkpoint chosen by the validation set loss.
2. Math reasoning: We finetune LLaMA2-7B (Touvron et al., 2023) on the MetaMathQA dataset (Yu et al., 2023), which contains 395K samples augmented from the training set of GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021).
We use test sets of GSM8K and MATH for evaluation and report results on the last checkpoint. For each test set, we report the exact match ratio against the ground truth.
3. Instruction following: We follow the experiment setup in Wu et al. (2024a) and Wu et al. (2024b) to finetune LLaMA2-7B with Ultrafeedback (Cui et al., 2023). Following previous works, we use Alpaca-Eval v1.0 (Li et al., 2023a) for evaluation and report win-rate against text-davinci-003 using GPT4-0613 as the annotator (OpenAI, 2024).
Implementation Details
We use the same hyperparameter configurations as Hu et al. (2023) without tuning for all methods. Details can be found in Appendix A.2. We denote this hyperparameter setup as LLM-Adapters. By default, we use a rank of 32. As math reasoning has a large training dataset, we set the rank to 64.